OpenStudy (anonymous):

Find f-1(x) (-1 is an inverse function). Sketch the graphs of y=f(x) and y=f-1(x) on the same coordinate axes. Answer for f(x)=2x+3 The answer I have is f-1(x)=x-3/2 Can someone just tell me if I was correct please?

OpenStudy (anonymous):

I have graphed the answer; however cannot figure exactly how to graph the other one.

OpenStudy (anonymous):

Take the original function y=2x+3. Solve for x, you have the inverse.

OpenStudy (anonymous):

y=2x+3 x=2y+3 x-3=2y x/2-3/2=y